OpenStudy (anonymous):

Solve for the values in [0,2pi] to make the following true. sin(5a) = -1 Thank you!

OpenStudy (anonymous):

explanation would be helpful :)

OpenStudy (anonymous):

Sin(5a)=-1 0<=a<=2pi LET 5a=x then sinx=-1 0<=x<=10pi

OpenStudy (anonymous):

NOw, solve for x

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

How/ where did you get 0<=x<=10pi??

OpenStudy (anonymous):

0<=a<=2pi 5*0<=5a<=5*2pi 0<=5a<=10pi Now, 5a=x So, 0<=x<=10pi

OpenStudy (anonymous):

and so.. with solving for x so.... x= -1/sin = -csc?

OpenStudy (anonymous):

OH I see that part better now.. thanks!!

OpenStudy (anonymous):

NOW, can u?

OpenStudy (anonymous):

ah. I'm confused about what to do from here on??

OpenStudy (anonymous):

is it something like sin (-1/sin) = -1? then you would get sin/sin = -1?

OpenStudy (anonymous):

sinx=-1 0<=x<=10pi 

OpenStudy (anonymous):

Now, use calculator

OpenStudy (anonymous):

I get -90? but how do I get the answer in pi format??

OpenStudy (anonymous):

ah. thank you for your help this late at night, I'm going to close this question to ask another question >__<